OpenStudy (anonymous):

The y-intercept of the line whose equation is 2x + 5y = 8 is 8??

OpenStudy (anonymous):

Set x at 0.

OpenStudy (anonymous):

y=8/5

OpenStudy (anonymous):

no,its y=8/5

OpenStudy (amistre64):

trivial but fun :) \(2x + 5y = 8\); /8 \(\cfrac{2x}{8} + \cfrac{5y}{8} = \cfrac{8}{8}\); simplify \(\cfrac{x}{4} + \cfrac{5y}{8} = 1\); just drop that "5" under \(\cfrac{x}{4} + \cfrac{y}{8/5} = 1\); just drop that "5" under read the intercepts from the bottoms

OpenStudy (amistre64):

did the latex mess that up on your screen?

OpenStudy (anonymous):

took a second or 2 to render and got "just drop that "5" under" twice, but it's fine.

OpenStudy (amistre64):

the "twice" but was typoed lol

OpenStudy (amistre64):

yes, that is a much nicer presentation :) you on opera or something?

OpenStudy (anonymous):

Firefox 5

OpenStudy (amistre64):

when I came back in it dropped down ... yay!!

OpenStudy (anonymous):

Odds and evens, I got no leader board.

OpenStudy (anonymous):

firefox has some built in math fonts and stuff and a nice extension called Firemath u would probably like.